COUNTING DISCRIMINANTS OF NUMBER FIELDS

HENRI COHEN FRANCISCO DIAZ Y DIAZ MICHEL OLIVIER

Abstract. This paper is intended as a summary of the conjectures and results obtained recently on the asymptotic and exact counting of relative and absolute extensions of number fields. No proofs are given.